A solution a prepared by dissolving 5.0g of salt in 550g of water what is the mass present of concentration of this solution
1
Expert's answer
2017-11-27T11:21:07-0500
w=(m(salt))/m(solution) =(m(salt))/(m(water)+m(salt))×100%=(5.0 g)/(550 g+5.0g)×100%=0.90%
Answer: 0.90%.